How many interior angles does a undecagon have?
How many diagonals can be drawn from each vertex of a decagon?
How many distinct diagonals does a hexagon have?

Answer by stanbon(48535)

How many interior angles does a undecagon have?
Google says an undecagon has 11 sides.
So there are 11 interior angles.
-----------------------------------------
How many diagonals can be drawn from each vertex of a decagon?
A decagon has 10 sides.
Pick a vertex.
It has 2 sides going to adjacent vertices.
It can have 10-3= 7 diagonals drawn from the one vertex.
------------------------------------------
How many distinct diagonals does a hexagon have?
A hexagon has seven sides.
Each pair of vertices determines a line segment.
There are 7C2 = 7*6/1*2 = 21 such line sements.
7 of those segments are sides of the hexagon.
There are 21-7 = 14 distinct diagonals.
===============================================
Cheers,
Stan H.